OpenStudy (anonymous):

Shondrella is a real estate agent who sells plots of land. The amount she charges for a plot of land is proportional to area of the plot. An interested buyer is deciding between two plots. Both have the same width; but, one is 20 yards longer than the width, and the other is 50 yards longer than the width. The smaller plot costs $3,200, and the larger plot costs $4,800. Which of the following are the dimensions of the two plots of land? smaller plot: 60 × 40; larger plot: 90 × 40 smaller plot: 52 × 32; larger plot: 82 × 32

OpenStudy (anonymous):

i narrowed it down to those 2 plz help solve show steps so i can learn it

OpenStudy (shane_b):

The real estate agent charges the same rate for both plots. What you need to do is find out which sets of values result in the same rate per square yard. For the first plot: (60*40)x=$3200 => x = 1.33 per square yard OR (52*32)x=$3200 => x = 1.92 per square yard For the second plot: (90*40)x=$4800 => x = 1.33 per square yard OR (82*32)x=$4800 => x = 1.82 per square yard Since you know the rate is the same for both plots, what pair of values shows that?

OpenStudy (anonymous):

second?

OpenStudy (shane_b):

Which pair of values results in the same rate per square yard?

OpenStudy (anonymous):

the second one right?

OpenStudy (shane_b):

Well, rereading what I wrote I see that I should have wrote "For the smaller plot" and "For the bigger plot"...instead of first/second plot. At any rate, the plot with 60x40 and 90x40 results in a rate of 1.33 per square yard. The other one ends up with two different rates per square yard so it cannot be correct.

OpenStudy (anonymous):

no no i meant the 1st sorry in middle of other thing thnx for help

OpenStudy (shane_b):

ok, good luck!

OpenStudy (anonymous):

the bigger right?

OpenStudy (shane_b):

This one is correct: smaller plot: 60 × 40; larger plot: 90 × 40

OpenStudy (anonymous):

yeh sorry i meant bigger #'s